Three capacitors are connected to each other as in alternatively between A to B of 4 micro F, 6 micro F and 8 micro F,to calculate equivalent capacitance from point A to B

1
Expert's answer
2021-06-07T09:32:26-0400

The reciprocal of the equivalent capacitance for capacitors linked in series is the total of the reciprocals of individual capacitances:

"\frac{1}{C_S}=\frac{1}{C_1}+\frac{1}{C_2}+\frac{1}{C_3}+\frac{1}{C_4}+.."

 a parallel combination of three capacitors, with one plate of each capacitor connected to one side of the circuit and the other plate linked to the other (a). Because the capacitors are parallel linked, they all have the same voltage V across their plates. Each capacitor in the parallel network, however, may hold a distinct charge.

"C_P=C_1+C_2+C_3+..."